OpenStudy (anonymous):

Carla is making ramps of different lengths. The ramps are made with rectangular and square wooden boards. The relationship between the length of the ramp in feet, f(n), and the number of square boards used, n, is represented by the function below. f(n) = 4n + 5 , where 20 ≤ n ≤ 30 Which statement is correct about the range of the length of the ramp for the given domain? It varies from 20 to 30. It varies from 20 to 125. It varies from 80 to 120. It varies from 85 to 125.

OpenStudy (jackellyn):

f(20)=4(20)+5 and f(30)=4(30)+5 will give you the range.

OpenStudy (anonymous):

It's D.

OpenStudy (anonymous):

yeah haha i did this problem last week

OpenStudy (anonymous):

lol